# Reshuffling Resampling Splits Can Improve Generalization of Hyperparameter Optimization

### v1

@@ 4.2 Experimental Results
In the following, we focus on the results obtained using ROC AUC. We present aggregated results over different tasks, learning algorithms and repetitions to get a general understanding of the effects. Unaggregated results and results involving accuracy and logloss can be found in Appendix F.
@@
In Figure 3 we exemplarily show how test performance develops over the course of an HPO run on a single task for different resamplings (with and without shuffling). Naturally, test performance does not necessarily increase in a monotonic fashion, and especially holdout without shuffling tends to be unstable. Its reshuffled version results in substantially better test performance.
@@
Next, we look at the relative improvement (compared to standard 5-fold CV, which we consider our baseline) with respect to test ROC AUC performance of the incumbent over time in Figure 4, i.e., the difference in test performance of the incumbent between standard 5-fold CV and a different resampling protocol; hence a positive difference tells us how much better in test error we are, if we would have chosen the other protocol instead 5-fold CV. We observe that reshuffling generally results in equal or better performance compared to the same resampling protocol without reshuffling. For 5-fold holdout and especially 5-fold CV and 5x 5-fold CV, reshuffling has a smaller effect on relative test performance improvement, as expected. Holdout is affected the most by reshuffling and results in substantially better relative test performance compared to standard HO. We also observe that an HPO protocol based on reshuffled holdout results in similar final test performance as standard 5-fold CV while overall being substantially cheaper due to requiring less model fits per HPC evaluation. In Appendix F.2, we further provide an ablation study on the number of folds when using $M$-fold holdout, where we observed that – in line with our theory – the more folds are used, the less reshuffling affects $M$-fold holdout.
Figure 4: Average improvement (compared to standard 5-fold CV) with respect to test performance (ROC AUC) of the incumbent over different tasks, learning algorithms and repetitions separately for increasing $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
However, this general trend can vary for certain combinations of classifier and performance metric, see the Appendix F. Especially for logloss, we observed that reshuffling rarely is beneficial; see the discussion in Section 5. Finally, the different resamplings generally behave as expected. The more we are willing to invest compute resources into a more intensive resampling like 5-fold CV or 5x 5-fold CV, the better the generalization performance of the final incumbent.
@@
Figure 5 shows that, generally, HEBO outperforms random search with respect to generalization performance (i.e., comparing both HEBO and random search under standard HO, or comparing both optimizers under reshuffled HO. More interestingly, both HEBO and random search strongly benefit from reshuffling. Moreover, the performance gap between HEBO and random search narrows when the resampling is reshuffled, which is an interesting finding of its own: As soon as we are concerned with generalization performance of HPO and not only investigate validation performance during optimization, the choice of optimizer might have less impact on final generalization performance compared to other choices such as whether the resampling is reshuffled during HPO or not. We present results for BO and reshuffling for different resamplings in Appendix F.
Figure 5: Average improvement (compared to random search on standard holdout) with respect to test performance (ROC AUC) of incumbent over tasks, learning algorithms and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@ 5 Discussion
To unveil the mechanisms underlying the reshuffling procedures, our theoretical analysis relied on an asymptotic approximation of the empirical loss surface. This was facilitated by a loss stability assumption regarding the learning algorithms that is generally rather mild; see the discussion in Bayle et al. (2020). However, it typically fails for highly sensitive losses, which has practical consequences. In fact, Figure 8 in Appendix F shows that reshuffling usually hurts generalization for the logloss and small sample sizes. It is still an open question whether this problem can be fixed by less naive implementations of the technique. Another limitation is our focus on generalization after search through a fixed, finite set of candidates. This largely ignores the dynamic nature of many HPO algorithms, which would greatly complicate our analysis. Finally, our experiments are limited in that we restricted ourselves to tabular data and binary classification and we avoided extremely small or large datasets.
@@
Several works suggest strategies to counteract this effect. First, LOOCVCV proposes a conservative choice of incumbents (Ng, 1997) at the cost of leave-one-out analysis or an additional hyperparameter. Second, it is possible to use an extra selection set (Igel, 2012; Lévesque, 2018; Mohr et al., 2018) at the cost of reduced training data, which was found to lead to reduced overall performance (Lévesque, 2018). Third, by using early stopping one can stop hyperparameter optimization before the generalization performance degrades again. This was so far demonstrated to be able to save compute budget at only marginally reduced performance, but also requires either a sensitivity hyperparameter or correct estimation of variance of the generalization estimate and was only developed for cross-validation so far (Makarova et al., 2022). Reshuffling itself is orthogonal to these proposals and a combination with the above-mentioned methods might result in further improvements.
@@
We further found, both theoretically and experimentally, that investing more resources in evaluating a single HPC can result in better final HPO performance. To reduce the computational burden on HPO again, we suggest further investigating the use of adaptive CV techniques, as proposed by Auto-WEKA (Thornton et al., 2013) or under the name Lazy Paired Hyperparameter Tuning (Zheng & Bilenko, 2013). Designing more advanced HPO algorithms exploiting the reshuffling effect should be a promising avenue for further research.
@@
Appendix A Extended Related Work
@@
In the last decade, most research on HPO has been concerned with constructing new algorithms that excel at finding configurations with a low estimated generalization error. Examples include BO variants such as as HEBO (Cowen-Rivers et al., 2022) or SMAC3 (Lindauer et al., 2022). Another direction of HPO research has been concerned with speeding up the HPO process to allow more efficient spending of compute resources. Multifidelity HPO, for example, turns the black box optimization problem into a gray box one by making use of lower fidelity approximations to the target function, i.e., using fewer numbers of epochs or subsets of the data for cheap low-fidelity evaluations that approximate the costly high-fidelity evaluation. Examples include bandit-based budget allocation algorithms such as Successive Halving (Jamieson & Talwalkar, 2016), Hyperband (Li et al., 2018) and their extensions that use non-random search mechanisms (Falkner et al., 2018; Awad et al., 2021; Mallik et al., 2023) or algorithms making use of multi-fidelity information in the context of BO (Swersky et al., 2014; Klein et al., 2017; Wu et al., 2020; Kadra et al., 2023).
@@
In general, little research has focused on the selection of resampling types, let alone the automated selection of resampling types (Guyon et al., 2010; Feurer et al., 2022). While we usually expect that a more intensive resampling will reduce the variance of the estimated generalization error and thereby improve the (rank) correlation between optimized validation and unbiased outer test performance within HPCs, this benefit is naturally offset by a higher computational expense. Overall, there is little research on which resampling method to use in practice for model selection, and we only know of a study for support vector machines (Wainer & Cawley, 2017), a simulation study for clinical prediction models (Dunias et al., 2024), a study on feature selection (Molinaro et al., 2005) and a study on fast CV (Bergman et al., 2024). The general consensus, that is in agreement with our findings, is that CV or repeated CV generally leads to better generalization performance. In addition, while there are theoretical works that compare the accuracy of estimating the generalization error of holdout and CV (Blum et al., 1999), our goals is to correctly identify a single solution, which generalizes well, see the excellent survey by Arlot & Celisse (2010) for a discussion on this topic.
For the design of machine learning experiments, Bouthillier et al. (2021) suggest to reshuffle the data prior to splitting it into the training, which is then used for HPO, and the test set. We followed their suggestion when designing our experiments and draw a new test sample for every replication, see Section 4.1 and Appendix E.
@@
A.1 Current Treatment of Resamplings in HPO Libraries and Software
@@
Appendix B Proofs of the Main Results
B.1 Proof of Theorem 2.1
@@ Appendix B Proofs of the Main Results
This assumption requires $g_{\bm{\lambda}}(\mathcal{T})$ to converge to some fixed prediction rule sufficiently fast and serves as a reasonable working condition for our purposes. It is satisfied, for example, when $\ell$ is the square loss and $g_{\bm{\lambda}}$ is an empirical risk minimizer over a hypothesis class $\mathcal{G}_{\bm{\lambda}}$ with finite VC-dimension. For further examples; see, e.g., Bousquet & Zhivotovskiy (2021), van Erven et al. (2015), and references therein. The assumption could be relaxed, but this would lead to a more complicated limiting distribution but with the same essential interpretation.
Theorem B.1.
@@
Appendix C Additional Result on the Density of Random HPC grids
Lemma C.1.
@@ Appendix C Additional Result on the Density of Random HPC grids
Appendix D Selected Validation Schemes
D.1 Definition of Index Sets
The following list provides formal definitions for the index sets $\mathcal{I}_{m,j}$ used in common validation schemes:
@@ D.1 Definition of Index Sets
D.2 Derivation of Reshuffling Parameters in Limiting Distribution
@@
For all schemes in the proposition, the probabilities are independent of the index $s$, so the average over $s=1,\dots,n$ can be omitted.
@@
    Hence, $\tau_{i,j,1}=1/\alpha$, and solving $\tau_{i,j,1}=\sigma^{2}\tau^{2}$ gives $\sigma^{2}=1/\alpha$, $\tau^{2}=1$.
@@
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times M/M=1/\alpha^{2}M^{2}.$
    Using $\alpha=1/M$, this implies $\sigma^{2}=1$, $\tau^{2}=1$.
@@
    For $i=j$, only $M$ probabilities in the double sum are non-zero, whence
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times M\times 1/M=1/\alpha^{2}M^{2}.$
@@
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times M^{2}\times 1/M^{2}=1/\alpha^{2}M^{2}.$
    Using $\alpha=1/M$, this implies $\sigma^{2}=1$, $\tau^{2}=1$.
@@
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times[M\times\alpha+(M-1)M\times\alpha^{2}]=1/\alpha M+(M-1)/M,$
    for all $i,j$, which implies $\sigma^{2}=1/M\alpha+(M-1)/M$, $\tau^{2}=1$.
@@
    This implies $\sigma^{2}=1/M\alpha+(M-1)/M$, $\tau^{2}=1/(1/M\alpha+(M-1)/M)$.
Remark D.1.
@@
Appendix E Details Regarding Benchmark Experiments
E.1 Datasets

### v2

@@ 4.2 Experimental Results
In the following, we focus on the results obtained using ROC AUC. We present aggregated results over different tasks, learning algorithms and replications to get a general understanding of the effects. Unaggregated results and results involving accuracy and logloss can be found in Appendix G.
@@
In Figure 3 we exemplarily show how test performance develops over the course of an HPO run on a single task for different resamplings (with and without reshuffling). Naturally, test performance does not necessarily increase in a monotonic fashion, and especially holdout without reshuffling tends to be unstable. Its reshuffled version results in substantially better test performance.
@@
Next, we look at the relative improvement (compared to standard 5-fold CV, which we consider our baseline) with respect to test ROC AUC performance of the incumbent over time in Figure 4, i.e., the difference in test performance of the incumbent between standard 5-fold CV and a different resampling protocol; hence a positive difference tells us how much better in test error we are, if we would have chosen the other protocol instead 5-fold CV. We observe that reshuffling generally results in equal or better performance compared to the same resampling protocol without reshuffling. For 5-fold holdout and especially 5-fold CV and 5x 5-fold CV, reshuffling has a smaller effect on relative test performance improvement, as expected. Holdout is affected the most by reshuffling and results in substantially better relative test performance compared to standard holdout. We also observe that an HPO protocol based on reshuffled holdout results in similar final test performance as standard 5-fold CV while overall being substantially cheaper due to requiring less model fits per HPC evaluation. In Appendix G.2, we further provide an ablation study on the number of folds when using $M$-fold holdout, where we observed that – in line with our theory – the more folds are used, the less reshuffling affects $M$-fold holdout.
Figure 4: Average improvement (compared to standard 5-fold CV) with respect to test performance (ROC AUC) of the incumbent over different tasks, learning algorithms and replications separately for increasing $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
However, this general trend can vary for certain combinations of classifier and performance metric, see Appendix G. Especially for logloss, we observed that reshuffling rarely is beneficial; see the discussion in Section 5. Finally, the different resamplings generally behave as expected. The more we are willing to invest compute resources into a more intensive resampling like 5-fold CV or 5x 5-fold CV, the better the generalization performance of the final incumbent.
@@
Figure 5 shows that, generally HEBO and SMAC3 outperform random search with respect to generalization performance (i.e., comparing HEBO and SMAC3 to random search under standard holdout, or comparing under reshuffled holdout). More interestingly, HEBO, SMAC3 and random search all strongly benefit from reshuffling. Moreover, the performance gap between HEBO and random search but also SMAC3 and random search narrows when the resampling is reshuffled, which is an interesting finding of its own: As soon as we are concerned with generalization performance of HPO and not only investigate validation performance during optimization, the choice of optimizer might have less impact on final generalization performance compared to other choices such as whether the resampling is reshuffled during HPO or not. We present results for BO and reshuffling for different resamplings in Appendix G.
Figure 5: Average improvement (compared to random search on standard holdout) with respect to test performance (ROC AUC) of the incumbent over tasks, learning algorithms and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@ 5 Discussion
To unveil the mechanisms underlying the reshuffling procedures, our theoretical analysis relies on an asymptotic approximation of the empirical loss surface. This allows us to operate on Gaussian loss surfaces, which exhibit convenient concentration and anti-concentration properties required in our proof. The latter are lacking for general distributions, which explains our asymptotic approach. The analysis was further facilitated by a loss stability assumption regarding the learning algorithms that is generally rather mild; see the discussion in Bayle et al. (2020). However, it typically fails for highly sensitive losses, which has practical consequences. In fact, Figure 9 in Appendix G shows that reshuffling usually hurts generalization for the logloss and small sample sizes. It is still an open question whether this problem can be fixed by less naive implementations of the technique. Another limitation is our focus on generalization after search through a fixed, finite set of candidates. This largely ignores the dynamic nature of many HPO algorithms, which would greatly complicate our analysis. Finally, our experiments are limited in that we restricted ourselves to tabular data and binary classification and we avoided extremely small or large datasets.
@@
Several works suggest strategies to counteract this effect. First, LOOCVCV proposes a conservative choice of incumbents (Ng, 1997) at the cost of leave-one-out analysis or an additional hyperparameter. Second, it is possible to use an extra selection set (Igel, 2012; Lévesque, 2018; Mohr et al., 2018) at the cost of reduced training data, which was found to lead to reduced overall performance (Lévesque, 2018). Third, by using early stopping one can stop hyperparameter optimization before the generalization performance degrades again. This was so far demonstrated to be able to save compute budget at only marginally reduced performance, but also requires either a sensitivity hyperparameter or correct estimation of the variance of the generalization estimate and was only developed for cross-validation so far (Makarova et al., 2022). Reshuffling itself is orthogonal to these proposals and a combination with the above-mentioned methods might result in further improvements.
@@
We further found, both theoretically and experimentally, that investing more resources when evaluating each HPC can result in better final HPO performance. To reduce the computational burden on HPO again, we suggest further investigating the use of adaptive CV techniques, as proposed by Auto-WEKA (Thornton et al., 2013) or under the name Lazy Paired Hyperparameter Tuning (Zheng & Bilenko, 2013). Designing more advanced HPO algorithms exploiting the reshuffling effect should be a promising avenue for further research.
@@
Appendix A Notation
Table 2: Notation table. We discuss all symbols used in the main paper.
$\bm{X}_{i}$ Random vector, describing the features
$Y_{i}$ Random variable, describing the target
$\bm{Z}_{i}=(\bm{X}_{i},Y_{i})$ Data point
$\mathcal{D}=\{\bm{Z}_{i}\}_{i=1}^{n}$ Dataset consisting of iid random variables
$n$ Number of observations
$g$ Inducer/ML algorithm
$h$ Model, created by the inducer via $h=g_{\bm{\lambda}}(\mathcal{D})$
$\bm{\lambda}$ Hyperparameter configuration
$\Lambda$ Finite set of all hyperparameter configurations
$J$ $|\Lambda|$, i.e., the number of hyperparameter configurations
$g_{\bm{\lambda}_{j}}$ Hyperparameterized inducer
$\mu(\bm{\lambda})$ Expected loss of a hyperparameterized inducer on the distribution of a dataset
$\ell({\bm{Z}},h)$ Loss of a model $h$ on a fresh observation $\bm{Z}$
$M$ Number of folds in M-fold cross-validation
$\alpha$ Percentage of samples to be used for validation
$\mathcal{I}_{1,j},\dots,\mathcal{I}_{M,j}\subset\{1,\dots,n\}$ $M$ sets of validation indices, to be used for evaluating $\bm{\lambda}_{j}$
$\mathcal{V}_{m,j}$ Validation data for fold $m$ and configuration $\bm{\lambda}_{j}$
$\mathcal{T}_{m,j}$ Training data for fold $m$ and configuration $\bm{\lambda}_{j}$
$L(\mathcal{V}_{m,j},g_{\bm{\lambda}_{j}}(\mathcal{T}_{m,j}))$ Validation loss for fold $m$ and configuration $\bm{\lambda}_{j}$
$\widehat{\mu}(\bm{\lambda}_{j})$ M-fold validation loss
$\sigma^{2}$ Increase in variance of validation loss caused by resampling
$\tau^{2}$ Decrease in correlation among validation losses caused by reshuffling
$\tau_{i,j,M}$ Resampling-related component of validation loss covariance
$K(\cdot,\cdot)$ Kernel capturing the covariance of the pointwise losses between two HPCs
$\epsilon(\bm{\lambda}_{j})$ Zero-mean Gaussian process, see Equation (2)
$d$ Number of hyperparameters
$\kappa$ Curvature constant of covariance kernel
$\eta$ Density of hyperparameter set $\Lambda$
$m$ Local curvature at the minimum of the loss surface $\mu$
$\underline{\sigma}$ Lower bound on the noise level
$B(\tau)$ Part of the regret bound penalizing reshuffling
$A(\tau)$ Part of the regret bound rewarding reshuffling
Appendix B Extended Related Work
@@
In the last decade, most research on HPO has been concerned with constructing new algorithms that excel at finding configurations with a low estimated generalization error. Examples include BO variants such as as HEBO (Cowen-Rivers et al., 2022) or SMAC3 (Lindauer et al., 2022). Another direction of HPO research has been concerned with speeding up the HPO process to allow more efficient spending of compute resources. Multifidelity HPO, for example, turns the black box optimization problem into a gray box one by making use of lower fidelity approximations to the target function, i.e., using fewer numbers of epochs or subsets of the data for cheap low-fidelity evaluations that approximate the costly high-fidelity evaluation. Examples include bandit-based budget allocation algorithms such as Successive Halving (Jamieson & Talwalkar, 2016), Hyperband (Li et al., 2018) and their extensions that use non-random search mechanisms (Falkner et al., 2018; Awad et al., 2021; Mallik et al., 2023) or algorithms making use of multi-fidelity information in the context of BO (Swersky et al., 2014; Klein et al., 2017; Wu et al., 2020; Kadra et al., 2023). Several works address the problem of speeding up cross-validation techniques and use techniques that could be described as grey box optimization techniques. Besides the ones mentioned in the main paper (Thornton et al., 2013; Zheng & Bilenko, 2013), it is possible to employ racing techniques for model selection in machine learning as demonstrated by Lang et al. (2015), and there has been a recent interest in methods that adapt the cost of running full cross-validation procedures (Bergman et al., 2024; Buczak et al., 2024).
@@
In general, little research has focused on the selection of resampling types, let alone the automated selection of resampling types (Guyon et al., 2010; Feurer et al., 2022). While we usually expect that a more intensive resampling will reduce the variance of the estimated generalization error and thereby improve the (rank) correlation between optimized validation and unbiased outer test performance within HPCs, this benefit is naturally offset by a higher computational expense. Overall, there is little research on which resampling method to use in practice for model selection, and we only know of a study for support vector machines (Wainer & Cawley, 2017), a simulation study for clinical prediction models (Dunias et al., 2024), a study on feature selection (Molinaro et al., 2005) and a study on fast CV (Bergman et al., 2024). In addition, ML-Plan (Mohr et al., 2018) proposed a two-stage procedure. In a first stage (search), the tool uses planning on hierarchical task networks to find promising machine learning pipelines on $70\%$ of the training data. In a second step (selection), it uses $100\%$ of the training data and retrains the most promising candidates from the search step. Finally, it uses a combination of the internal generalization error estimation that was used during search and the $0.75$ percentile of the generalization error estimation from the selection step to make a more unbiased selection of the final model. The paper found that this improves performance over using only regular cross-validation for search and selection. The general consensus, that is in agreement with our findings, is that CV or repeated CV generally leads to better generalization performance. In addition, while there are theoretical works that compare the accuracy of estimating the generalization error of holdout and CV (Blum et al., 1999), our goals is to correctly identify a single solution, which generalizes well, see the excellent survey by Arlot & Celisse (2010) for a discussion on this topic.
Bouthillier et al. (2021) studied the sources of variance in machine learning experiments, and find that the split into training and test data has the largest impact. Consequently, they suggest to reshuffle the data prior to splitting it into the training, which is then used for HPO, and the test set. We followed their suggestion when designing our experiments and draw a new test sample for every replication, see Section 4.1 and Appendix F. This dependence on the exact split was further already discussed in the context of how much the outcome of a statistical test on results of machine learning experiments depended on the exact train-test split (Bouckaert, 2004).
@@
B.1 Current Treatment of Resamplings in HPO Libraries and Software
@@
Appendix C Proofs of the Main Results
C.1 Proof of Theorem 2.1
@@ Appendix C Proofs of the Main Results
This assumption requires $g_{\bm{\lambda}}(\mathcal{T})$ to converge to some fixed prediction rule sufficiently fast and serves as a reasonable working condition for our purposes. It is satisfied, for example, when $\ell$ is the square loss and $g_{\bm{\lambda}}$ is an empirical risk minimizer over a hypothesis class $\mathcal{G}_{\bm{\lambda}}$ with finite VC-dimension. For further examples, see, e.g., Bousquet & Zhivotovskiy (2021), van Erven et al. (2015), and references therein. The assumption could be relaxed, but this would lead to a more complicated limiting distribution but with the same essential interpretation.
Theorem C.1.
@@
Appendix D Additional Results on the Density of Random HPC Grids
Lemma D.1.
@@ Appendix D Additional Results on the Density of Random HPC Grids
Appendix E Selected Validation Schemes
E.1 Definition of Index Sets
Recall:
@@ E.1 Definition of Index Sets
E.2 Derivation of Reshuffling Parameters in Limiting Distribution
@@
For all schemes in the proposition, the probabilities are independent of the index $s$, so the average over $s=1,\dots,n$ can be omitted. We now verify the constants $\sigma,\tau$ from Table 1.
@@
    Hence,
    $\tau_{i,j,1}=1/\alpha=1/\alpha\times 1=\sigma^{2}\times\tau^{2}.$
@@
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times M/M=1/\alpha^{2}M^{2}=1\times 1=\sigma^{2}\times\tau^{2},$
    where we used $\alpha=1/M$.
@@
    For $i=j$, only $M$ probabilities in the double sum are non-zero. Also using $\alpha=1/M$, we get
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times M\times 1/M=1=\sigma^{2}.$
@@
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times M^{2}\times 1/M^{2}=1\times 1=\sigma^{2}\times\tau^{2}.$
@@
    $\tau_{i,j,M}=\frac{1}{M^{2}\alpha^{2}}\times[M\times\alpha+(M-1)M\times\alpha^{2}]=[1/\alpha M+(M-1)/M]\times 1=\sigma^{2}\times\tau^{2}.$
    for all $i,j$.
@@
    This implies that (1) holds with $\sigma^{2}=1/M\alpha+(M-1)/M$, $\tau^{2}=1/(1/M\alpha+(M-1)/M)$.
Remark E.1.
@@
Appendix F Details Regarding Benchmark Experiments
F.1 Datasets